# Expressivity of Representation Learning on Continuous-Time Dynamic Graphs: An Information-Flow Centric Review

### v1

@@ 2.1 Key Components of CTDG Representation Learning
    $s_{u}(t)=\textsc{MemUpd}([s_{u}(t^{-}),s_{v}(t^{-}),t-t^{-},e_{u,v}^{t}])$ (1)
    $s_{v}(t)=\textsc{MemUpd}([s_{v}(t^{-}),s_{u}(t^{-}),t-t^{-},e_{u,v}^{t}])$ (2)
@@ 2.2 Expressivity Framework Based on Information Flow (IF)
Specifically, our goal is to examine the change in node embeddings between different consecutive time points to understand how events impact the node representations. Naturally and intuitively, the nodes directly involved in the events are expected to undergo the most significant changes in their embeddings, while the representations of other nodes are less affected. This analysis provides valuable insights into how the graph’s structure evolves in response to events, offering a detailed view of how different parts of the graph react. By studying the evolution of node embeddings, we can also assess how various methods handle information propagation across the graph. This sheds light on the strengths and limitations of different techniques, offering guidance on their applicability to specific use cases where certain types of event-driven changes are more prevalent.
@@ 3.1 Method Categorization Inspired by Theoretical Framework
We start with reviewing various GRL methods for CTDGs, categorized based on the theoretical insights outlined in Section 2.2. The expressivity framework provides a systematic basis for evaluating the strengths and limitations of each method across different graph types. At a high level (cf. Figure 1), the categorization distinguishes between methods that use explicit MP mechanisms and those that rely on direct representation updates driven by events. Additionally, we examine strategies for memory aggregation and updates, emphasizing their influence on model performance, expressivity, and computational complexity.
@@ 3.2.1 Pretext tasks for pretraining on CTDG
However, contrastive pretraining has been less explored in CTDGs. Only a few studies (Chen et al., 2023; Sun et al., 2022) have extended this technique to DG. In particular, DySubC (Chen et al., 2023) introduced a contrastive learning framework designed to maximize the mutual information between each node’s embedding and the representation of its associated temporal subgraph. DySubC consists of generating time-aware subgraphs for each node by capturing both structural information from the node’s neighborhood and temporal information from edge timestamps, thereby encoding the dynamics of interactions over time. It then constructs positive and negative samples: positive pairs consist of the central node and its time-weighted subgraph representation, while negative samples are created by applying structural and temporal perturbations, such as shuffling the subgraph structure or removing temporal weight edges.
@@ 4 Empirical Validation
This section presents an empirical study to validate the theoretical insights and categorization proposed earlier. Specifically, we evaluate the effectiveness of reviewed methods and identify the types of graphs where they perform best. We first outline the experimental setup, followed by a discussion of the empirical findings. Additionally, an analysis of the tightness of the proposed upper bounds is provided in Appendix E.
@@ 4.1 Experimental Setup
Our experimental focus is on the popular link prediction task, chosen as a representative challenge within the broader scope of CTDGs. We examine a diverse selection of models from those reviewed in Section 3, focusing on models that capture different facets of CTDG methodologies: (I) JODIE (Kumar et al., 2019a), which models evolving embeddings over time; (II) DyRep (Trivedi et al., 2019), designed for capturing dynamic node interactions; (III) TGN (Rossi et al., 2020b), i.e., Temporal Graph Networks for inductive learning; (IV) TGAT (Xu et al., 2020), known for its attention-based approach for handling temporal information; (V) CTAN (Gravina et al., 2024), focused on modeling long-range dependencies within graphs. Experimental details and hyper-parameters used to train/test the model are provided in Appendix F.2.
@@
In alignment with the selected methods, we choose a set of both synthetic and real-world DG datasets which are representative to the main application areas and unique characteristics of CTDGs. Additional details on each dataset, including design considerations and properties, are provided in Appendix F.1.
@@
Long-range graphs are those requiring long-range reasoning to achieve strong performance, as distant nodes or graph regions can influence each other. To capture these interactions, specific information propagation mechanisms are necessary, as conventional MP often fails to bridge these long-range dependencies effectively. Evaluating performance on these graphs is crucial to assess a method’s ability to transfer information across distant parts of the graph. For this setting, we use an approach similar to the one proposed by Gravina et al. (2024) where we consider a temporal version of the PascalVOC-SP graph (Dwivedi et al., 2022). Additional information about the generation is provided in Appendix F.1.
@@
Bi-partite graphs play a crucial role in applications like recommender systems, a prominent use case for CTDGs. To evaluate how different methods handle interactions between disjoint node sets, which are critical for recommendation quality, we use the TGBL-Wiki dataset from the TGB benchmarks (Huang et al., 2024). This dataset represents a bi-partite network where nodes are editors and Wiki pages, with edges added when an editor modifies a page at a specific timestamp. Edges carry text features derived from the page edits. Dataset statistics can be found in Table 5 in Appendix F.1.
@@
Community-based graphs consist of nodes grouped into distinct clusters, with relatively few inter-community links compared to intra-community connections. A Barbell graph is a representative example of this structure, where inter-community links act as bottlenecks (Alon & Yahav, 2021), making these graphs challenging for temporal MP schemes due to limited information propagation between communities. To evaluate how well a benchmark captures intra-community links, we generate a synthetic dataset using the Stochastic Block Model (SBM) with the NetworkX (Hagberg et al., 2008) package. This involves creating $B$ dense communities with connections randomly sampled based on a normalized density over a time horizon. As time progresses, the graph evolves, becoming increasingly connected across clusters. For evaluation, we leverage the connection sampling densities between clusters to rank nodes across active communities, measuring performance using Normalized Discounted Cumulative Gain (NDCG) (Järvelin & Kekäläinen, 2002). More details on the parameters, generation and evaluation process are provided in Table 4 in Appendix F.1.
@@ 4.2 Results Analysis
Table 3 summarizes the average test and validation MRR (Mean Reciprocal Rank) scores, along with standard deviations, for various methods on the real-world bi-partite dataset TGBL-Wiki. The results show that JODIE performs similarly to the MP based methods DyRep and TGAT. While JODIE might theoretically be expected to excel on bi-partite graphs, the dataset’s low “surprise factor” — a measure of how predictable the interactions between nodes are — limits its advantages. In this setting, capturing long-range temporal and structural dependencies becomes more critical for the downstream task. Consequently, methods like CTAN and TGN, which are better equipped to model such dependencies, achieve significantly superior performance.
@@ Appendix A Preliminaries
Problem Set-up For our theoretical analysis, we consider that all the models are using $1-Lipschitz$ continuous activation functions. We note that this is the case for the majority of classically used functions (ReLU, LeakyReLU, tanh, and others) (Virmaux & Scaman, 2018). As our goal is to understand the effect of each component on the model’s evolving dynamics, we consider the most general use-case of a CTDG function. Typically, we consider that our model $f$ follows the proposed components in Section 2.1. Specifically, we consider the model to based on the described temporal message-passing (either through a GCN-like aggregation or attention-based). We also consider that it makes use of the memory state and not only the node’s individual representation. We additionally consider that the model encompasses a Linear temporal projection function to take into account temporal inactivity.
@@ B.1 For nodes not involved in the event and for which $L<d(u,i)$
In this part, we consider the nodes that are not directly connected in the events and that are distant from the nodes involved in the events. Specifically, we consider that the number of propagation $L<d(u,i)$ with $u$ being the node and $i$ being the node included in the event and $d$.
In this proof, we consider that we are dealing with a GCN-like propagation:
@@
$=\left\lVert\sigma^{(l)}\!\left(\sum_{v\in\mathcal{N}(u)\bigcup\{u\}}\frac{W^{(\ell)}h_{v,t+1}^{(\ell-1)}}{\sqrt{(1+\text{deg(u)})(1+\text{deg(v)})}}\right)-\sigma^{(l)}\!\left(\sum_{v\in\mathcal{N}(u)\bigcup\{u\}}\frac{W^{(\ell)}h_{v,t}^{(\ell-1)}}{\sqrt{(1+\text{deg(u)})(1+\text{deg(v)})}}\right)\right\lVert$
@@
By iteration, we get the following:
@@ C.2 For any node $u$ not involved in the event and for which $L\geq d(u,i)$
Appendix D Specific Cases
D.1 In the case of MLPs
A typical Multi-Layer Perceptron (MLP) with $L$ layers implements the following:
$h_{1}$ $=\phi(W_{1}x+b_{1})$
$h_{2}$ $=\phi(W_{2}h_{1}+b_{2})$
[…]
$h_{L}$ $=\phi(W_{L}h_{L-1}+b_{L})$
where $\phi$ is an 1-Lipschitz activation function applied element-wise. Then:
$L_{\text{MLP}}\leq\prod_{l=1}^{L}\rVert W_{l}\lVert_{2}$ (12)
Hence, for methods using MLPs, we have the following Lipschitz bound:
$\lVert h_{L}(x_{1})-h_{L}(x_{2})\rVert\leq L_{\text{MLP}}\lVert x_{1}-x_{2}\rVert.$ (13)
Appendix E On the Tightness of the Provided Upper-Bound
@@
with $w_{2}$ being the learned weight matrix used with the edge features and ${e^{\prime}}_{i,j}=[e_{i,j}\mid\mid\Delta_{t}]$ is the concatenated vector of the edge features and the timestamp as explained in Rossi et al. (2020b). By following this formulation and using the triangular inequality, we can adapt the upper-bound provided in 10 as follows:
@@
Appendix F Datasets and Implementation Details
F.1 Synthetic Datasets

### v2

@@ 2.1 Key Components of CTDG Representation Learning
    $s_{u}(t)=\textsc{MemUpd}([s_{u}(t^{-}),s_{v}(t^{-}),t-t^{-},e_{u,v}^{t}]);\hskip 1.00006pts_{v}(t)=\textsc{MemUpd}([s_{v}(t^{-}),s_{u}(t^{-}),t-t^{-},e_{u,v}^{t}])$ (1)
@@ 2.2 Expressivity Framework Based on Information Flow (IF)
Specifically, our goal is to examine the change in node embeddings between consecutive time points to understand how events impact the node representations. Naturally and intuitively, the nodes directly involved in the events are expected to undergo the most significant changes in their embeddings, while the representations of other nodes are less affected. This analysis provides valuable insights into how the graph’s structure evolves in response to events, offering a detailed view of how different parts of the graph react. By studying the evolution of node embeddings, we can also assess how various methods handle information propagation across the graph. This sheds light on the strengths and limitations of different techniques, offering guidance on their applicability to specific use cases where certain types of event-driven changes are more prevalent.
@@ 3.1 Method Categorization Inspired by Theoretical Framework
The proposed expressivity framework in Section 2.2 provides a systematic basis for evaluating the strengths and limitations of each method across different graph types. Typically, we can see that a model’s expressivity depends on the message-passing framework and the usage of memory where relying only on one single-hop update with minimal memory might have a small upper-bound on its IF for distance node, limiting therefore its capacity to melodize the event occurrence. Based on these insights, we introduce a categorization (Figure 1) to distinguish between the different choices. An additional visual summary is provided in Appendix F.
@@ 3.2.1 Pretext tasks for pretraining on CTDG
However, contrastive pretraining has been less explored in CTDGs. Only a few studies (Chen et al., 2023; Sun et al., 2022) have extended this technique to DG. In particular, DySubC (Chen et al., 2023) introduced a contrastive learning framework designed to maximize the mutual information between each node’s embedding and the representation of its associated temporal subgraph. DySubC consists of generating time-aware subgraphs for each node by capturing both structural information from the node’s neighborhood and temporal information from edge timestamps, thereby encoding the dynamics of interactions over time. It then constructs positive and negative samples: positive pairs consist of the central node and its time-weighted subgraph representation, while negative samples are created by applying structural and temporal perturbations, such as shuffling the subgraph structure or removing temporal weight edges. In the same line, CLDG (Xu et al., 2023) approaches contrastive learning by leveraging temporal translation invariance through a sampling layer that extracts persistent signals across snapshots. The aim is to maintain a node’s local and global representations while mitigating semantic shifts introduced by temporal augmentations. Additionally, IDOL makes use of a specialized Personalized PageRank mechanism to capture topological changes, and integrates a topology-monitorable sampling strategy yielding better representations.
@@ 4 Empirical Validation
This section presents an empirical study to validate the theoretical insights and categorization proposed earlier. Specifically, we evaluate the effectiveness of reviewed methods and identify the types of graphs where they perform best. We first outline the experimental setup, followed by a discussion of the empirical findings. Additionally, an analysis of the tightness of the proposed upper bounds is provided in Appendix D.
@@ 4.1 Experimental Setup
Our experimental focus is on the popular link prediction task, chosen as a representative challenge within the broader scope of CTDGs. We examine a diverse selection of models from those reviewed in Section 3, focusing on models that capture different facets of CTDG methodologies: (I) JODIE (Kumar et al., 2019a), which models evolving embeddings over time; (II) DyRep (Trivedi et al., 2019), designed for capturing dynamic node interactions; (III) TGN (Rossi et al., 2020b), i.e., Temporal Graph Networks for inductive learning; (IV) TGAT (Xu et al., 2020), known for its attention-based approach for handling temporal information; (V) CTAN (Gravina et al., 2024), focused on modeling long-range dependencies within graphs. Experimental details and hyper-parameters used to train/test the model are provided in Appendix G.2.
@@
In alignment with the selected methods, we choose a set of both synthetic and real-world DG datasets which are representative to the main application areas and unique characteristics of CTDGs. Additional details on each dataset, including design considerations and properties, are provided in Appendix G.1.
@@
Long-range graphs are those requiring long-range reasoning to achieve strong performance, as distant nodes or graph regions can influence each other. To capture these interactions, specific information propagation mechanisms are necessary, as conventional MP often fails to bridge these long-range dependencies effectively. Evaluating performance on these graphs is crucial to assess a method’s ability to transfer information across distant parts of the graph. For this setting, we use an approach similar to the one proposed by Gravina et al. (2024) where we consider a temporal version of the PascalVOC-SP graph (Dwivedi et al., 2022). Additional information about the generation is provided in Appendix G.1.
@@
Bi-partite graphs play a crucial role in applications like recommender systems, a prominent use case for CTDGs. To evaluate how different methods handle interactions between disjoint node sets, which are critical for recommendation quality, we use the TGBL-Wiki dataset from the TGB benchmarks (Huang et al., 2024). This dataset represents a bi-partite network where nodes are editors and Wiki pages, with edges added when an editor modifies a page at a specific timestamp. Edges carry text features derived from the page edits. Dataset statistics can be found in Table 6 in Appendix G.1.
@@
Community-based graphs consist of nodes grouped into distinct clusters, with relatively few inter-community links compared to intra-community connections. A Barbell graph is a representative example of this structure, where inter-community links act as bottlenecks (Alon & Yahav, 2021), making these graphs challenging for temporal MP schemes due to limited information propagation between communities. To evaluate how well a benchmark captures intra-community links, we generate a synthetic dataset using the Stochastic Block Model (SBM) with the NetworkX (Hagberg et al., 2008) package. This involves creating $B$ dense communities with connections randomly sampled based on a normalized density over a time horizon. As time progresses, the graph evolves, becoming increasingly connected across clusters. For evaluation, we leverage the connection sampling densities between clusters to rank nodes across active communities, measuring performance using Normalized Discounted Cumulative Gain (NDCG) (Järvelin & Kekäläinen, 2002). More details on the parameters, generation and evaluation process are provided in Table 5 in Appendix G.1.
@@ 4.2 Results Analysis
Table 3 summarizes the average test and validation MRR (Mean Reciprocal Rank) scores, along with standard deviations, for various methods on the real-world bi-partite dataset TGBL-Wiki. The results show that JODIE performs similarly to the MP based methods DyRep and TGAT. While JODIE might theoretically be expected to excel on bi-partite graphs, the dataset’s low “surprise factor” — a measure of how predictable the interactions between nodes are — limits its advantages. In this setting, capturing long-range temporal and structural dependencies becomes more critical for the downstream task. Consequently, methods like CTAN and TGN, which are better equipped to model such dependencies, achieve superior performance.
@@ 5 Conclusion
Acknowledgments
We would like to thank the entire ACE (AI Center of Excellence) at King (a part of Microsoft) for their continuous support, insightful discussions, and collaborative spirit throughout this project.
We express our sincere gratitude to King’s OSC (Open Source Council) for their invaluable support throughout the open source release process. Special thanks go to Colin Cashin, Nate Cleveland, Hugo Cura, Paul Jordaan, David Lorenzo, and Raul Pareja for their dedicated efforts in facilitating the publication and code release process. We are also deeply appreciative of King’s CTO, Eric Bowman, for his careful review and approval of our code (https://github.com/king/ctdg-info-flow).
Additionally, we are grateful for the diligent review of our manuscript from legal and communications perspectives. Special thanks to Alexandra Stark and Tim Elgar for their essential contributions in ensuring alignment with legal and communications standards.
This work was partially funded by the Wallenberg AI, Autonomous Systems and Software Program (WASP). We also acknowledge the collaboration between King/Microsoft and KTH Royal Institute of Technology, which made this research possible.
@@ Appendix A Preliminaries
Problem Set-up. As our goal is to understand the effect of each component on the model’s evolving dynamics, we consider the most general use-case of a CTDG function. Typically, we consider that our model $f$ follows the proposed components in Section 2.1. Specifically, we consider the model to based on the described temporal message-passing (either through a GCN-like aggregation or attention-based). We also consider that it makes use of the memory state and not only the node’s individual representation. We additionally consider that the model encompasses a Linear temporal projection function to take into account temporal inactivity.
Theoretical Assumptions. For our theoretical analysis, we assume that all models use $1$-Lipschitz continuous activation functions, which covers the majority of commonly used functions, including ReLU, LeakyReLU, and TanH (Virmaux & Scaman, 2018). We also focus on the effect of a single event, in contrast to the more practical setting where a batch of events is considered. Under this single-event assumption, the node distribution can be treated as static between two relevant time shot of the graph at $t$ and $t+1$. While this assumption may not strictly hold when processing batches, we consider that deriving the specific bounds will yield the same theoretical insights on the considered method’s ability to propagate information.
@@
Proof’s Intuition. In our proof, we want to see the effect of an event on the node representation of a node $u$ within our graph. Typically, the set of nodes can be divided into two main categories. One first set which is within a reachable distance of where the event happened in respect to the number of message-passing layers. A second set, which is not reached based on the chosen number of message-passing. We approached the proof therefore by considered each set independently, where we upper-bound the expected change in each set. This latter division gives an intuition on how the considered model $f$ interacts with the different sets and consequently reflects its ability to propagate information within the graph after an event occurrence.
@@ B.1 For nodes not involved in the event and for which $L<d(u,i)$
In this part, we consider the nodes that are not directly connected in the events and that are distant from the nodes involved in the events. Specifically, we consider that the number of propagation $L<d(u,i)$ with $u$ being the node and $i$ being the node included in the event and $d$. In this proof, we consider that we are dealing with a GCN-like propagation, we can therefore right the following:
@@
$=\lVert\sigma^{(l)}\!\left(\sum_{v\in\mathcal{N}(u)\bigcup\{u\}}\frac{W^{(\ell)}h_{v,t+1}^{(\ell-1)}}{\sqrt{(1+\text{deg(u)})(1+\text{deg(v)})}}\right)$
$\hskip 155.00023pt-\sigma^{(l)}\!\left(\sum_{v\in\mathcal{N}(u)\bigcup\{u\}}\frac{W^{(\ell)}h_{v,t}^{(\ell-1)}}{\sqrt{(1+\text{deg(u)})(1+\text{deg(v)})}}\right)\lVert$
@@
We see that the right term on the resulting inequality is related to the one we are considering. Therefore, by iteratively doing the same process, we get the following:
@@ B.2 For nodes not involved in the event and for which $L\geq d(u,i)$
Proof’s Intuition. Similar to the previous proof, we consider in this case that the model is rather based on an attention framework to aggregate the information within the neighborhood. We start therefore by analyzing the effect of this addition. We afterwards, and in the same way as the GCN-case, derived the upper-bound for each set of nodes based on the distance to the event.
@@ C.2 For any node $u$ not involved in the event and for which $L\geq d(u,i)$
Appendix D On the Tightness of the Provided Upper-Bound
@@
with $w_{2}$ being the learned weight matrix used with the edge features and ${e^{\prime}}_{i,j}=[e_{i,j}\mid\mid\Delta_{t}]$ is the concatenated vector of the edge features and the timestamp as explained in Rossi et al. (2020b). By following this formulation and using the triangular inequality, we can adapt the upper-bound provided in 9 as follows:
@@
Appendix E Empirical Insights of the Information-Flow
As discussed in Section 2.2, the concept of Information-Flow serves as a metric for assessing the influence of an event on the graph’s node embeddings. In particular, when a new edge is added (i.e., an event occurs), it is expected that the node embeddings will be updated accordingly. If no update is observed, for example, in an edge prediction scenario, the corresponding probabilities remain unchanged, indicating that the event provides no additional information. Ideally, the embeddings of all nodes in the graph should be updated in response to the event.
From the theoretical results, one key determinant of improved Information-Flow is the number of message-passing layers. To illustrate this, the normalized average norm difference of node embeddings is evaluated following an event in the TGBL-Wiki dataset, while varying the number of message-passing layers. Specifically, the TGAT model is considered with different configurations ($L=1,2,3$) during inference. As depicted in Figure 3 (left axis), models with a larger number of layers exhibit a greater degree of information propagation across the graph.
Figure 3: Information Flow change after an event occurring and the Average Test MRR on the TGBL-Wiki dataset for different number of layers.
However, it should be noted that although increased information flow is generally desirable, it does not necessarily lead to superior downstream performance. This discrepancy is often attributed to over-smoothing, a phenomenon well-documented in the literature. In certain graph topologies, such as bipartite structures, a model with a smaller number of layers can outperform deeper models. Figure 3 (right axis) further reports the Average Test MRR for the same TGAT models, highlighting this effect.
Appendix F Visual Summary of the Theoretical Insights
Our work position itself as a review of the available methods in the context of CTDGs. Specifically, our primary goal is to offer a unifying perspective on CTDG methods, especially in self-supervised settings, so that given a graph structure and constraints, the user can choose the right method to be used to maximize the performance. In this perspective, our provided "Information-Flow" in Section 2.2 studies how a method’s components can affect its ability to propagate and update information after an event occurrence. This is relevant to understand how different methods shall interact with different topologies. In this perspective, Table 4 provides a visual summary of the different insights that were derived from our theoretical analysis and the computed upper bounds. Specifically, it highlights the key components influencing the “expressivity” of a temporal graph function, detailing their functional roles, associated mathematical constraints, and practical implications. The table summarize how factors such as memory usage, projection mechanisms, neighborhood depth, event features, and attention mechanisms contribute to the propagation of information over time, while also delineating the trade-offs in computational complexity and model stability.
Component Description Bound Terms Practical Effect
Memory Usage Tracks historical interactions for each node. $\Delta(s_{i})$: Change in memory state for nodes involved in events. Improves temporal dependency modeling; sensitive to memory size and update mechanism.
Temporal Projection Projects node representations forward during inactivity. $|W_{t}|$: Temporal projection weight matrix norm. Captures time-varying node states but adds sensitivity to temporal sparsity.
Aggregator Type Mechanism for combining neighborhood information (e.g., summation, attention). Aggregator-dependent: $\prod_{l=1}^{L}|W^{(l)}|$ for GCN, degree-based terms for attention. Influences expressivity and computational cost; enables fine-grained updates.
Neighborhood Depth ($L$) Number of layers in the message-passing scheme, affecting the propagation range. $\hat{w}_{u},\hat{w}_{j}$: Normalized walks or shortest paths. Controls the range of information flow but risks over-smoothing or over-squashing.
Event Features Includes edge-related attributes (e.g., interaction type, timestamps). $|W_{t}|+\mathcal{B}$: Interaction with edge features through projection and event-related adjustments. Enhances representation with context-specific interactions but may increase model complexity.
Attention Mechanism Dynamically weighs neighbor contributions using learnable parameters. Degree-based terms ($deg(u)$) and learned weights ($W$). Enables selective information propagation, particularly useful for graphs with heterogeneous links.
Graph Topology Structural characteristics such as sparsity, degree distribution, and community structures. $\hat{w}_{u},\hat{w}_{j}$: Impact of topology on walk-based terms. Affects long-range dependency modeling and suitability for specific graph types (e.g., bipartite).
Table 4: Visual Representation of the theoretical insights from the computer upper-bounds in Theorem 2.3 and 2.2.
Appendix G Datasets and Implementation Details
G.1 Synthetic Datasets